# SlowFast-LLaVA: A strong training-free baseline for video large language models

### v1

@@ 5 Conclusion
We present SF-LLaVA, a training-free Video LLM that is built upon LLaVA-NeXT and requires no additional fine-tuning to work effectively for various video tasks. Especially, we propose a SlowFast design that uses two-stream inputs for Video LLMs. It aggregates frame features as an effective video representation that can capture both detailed spatial semantics and long-range temporal context. Our experiments on a diverse set of 8 video benchmarks demonstrate the superiority of SF-LLaVA, where it outperforms existing training-free methods. On some benchmarks, SF-LLaVA achieves on-par or even better results than state-of-the-art SFT Video LLMs that have been extensively fine-tuned on large-scale video data. We hope SF-LLaVA can serve as a simple but strong baseline in the whole picture of Video LLMs, and our ablation on its design choices can provide valuable insights for future research on modeling video representation for Multimodal LLMs.

### v2

@@ 5 Conclusion
We present SF-LLaVA, a training-free Video LLM that is built upon LLaVA-NeXT and requires no additional fine-tuning to work effectively for various video tasks. Especially, we propose a SlowFast design that uses two-stream inputs for Video LLMs. It aggregates frame features as an effective video representation that can capture both detailed spatial semantics and long-range temporal context. Our experiments on a diverse set of 8 video benchmarks demonstrate the effectiveness of SF-LLaVA, where it outperforms existing training-free methods. On some benchmarks, SF-LLaVA achieves on-par or even better results than state-of-the-art SFT Video LLMs that have been extensively fine-tuned on large-scale video data. We hope SF-LLaVA can serve as a simple but strong baseline in the whole picture of Video LLMs, and our ablation on its design choices can provide valuable insights for future research on modeling video representations for Multimodal LLMs.